OpenStudy (anonymous):

calculate dy/dx as a function of theta. determine the polar coordinates (r,theta), of any horizontal and vertical tangents, and convert them to Cartesian coordinates. for the fucntion r(theta)=(4/(3-sin(theta)))

OpenStudy (anonymous):

\(\large \frac{dy}{dx}=\frac{dy/d\theta}{dx/d\theta} \) where x=rcos(theta) and y=rsin(theta) since \(\large r(\theta)=\frac{4}{3-sin\theta} \), then \(\large x=\frac{4cos\theta}{3-sin\theta} \), and \(\large y=\frac{4sin\theta}{3-sin\theta} \) can you calculate dy/d(theta) and dx/d(theta) from here?

OpenStudy (anonymous):

Im struggling to solve for dy/d(theta) and dx/d(theta) from here. would you use the quotient rule on to find the derivatives for each x and y functions?